OpenStudy (anonymous):

what is the solution of 4=(8/x+20)

OpenStudy (raden):

4=(8/x) + 20 or 4 = 8/(x+20) ??

OpenStudy (anonymous):

x=-1/2

OpenStudy (anonymous):

then second one

OpenStudy (anonymous):

You can do cross-multiply of x+20 times 4. That solution could then be put equal to 8, and now it's simpler to solve.

OpenStudy (anonymous):

all i need to know for this one is what are the numbers that cancle out

OpenStudy (anonymous):

Remember when you're doing the cross multiplication, you are using distribution. So 4(x+20)=8.

OpenStudy (raden):

divided by 4 from both sides,